OpenStudy (anonymous):

which of the following is the simplified form of (x+4)/(2x+6) + (3)/(x^2-9) a) x+2/2(x+3) b) x+2/2(x-3) c)x-2/2(x-3) d)x-2/2(x+3)

OpenStudy (anonymous):

[(x+4)/2(x+3)]+[3/(x-3)(x+3)] =[(x+4)(x-3)+3*2]/2(x-3)(x+3) =x^2+x-12+6/2(x-3)(x+3) =(x-3)(x+2)/2(x-3)(x+3) =(x+2)/2(x+3)

OpenStudy (anonymous):

(x+4)/(2x+6) + (3)/(x^2-9) =(x+4)(x-3)/2(x+3)(x-3) + (3)(2)/2(x+3)(x-3) =((x+4)(x-3)+6)/2(x+3)(x-3) =(x^2-x-6)/2(x+3)(x-3) =(x-3)(x+2)/2(x+3)(x-3) =(x+2)/2(x+3)

OpenStudy (anonymous):

thanks you guys!